You need 0.56 mol of electrons to deposit a thin covering of silver on a part that you are using for a prototype machine. How long would you need to run a current of 5.0 A to obtain this number of moles?


1
Expert's answer
2021-12-15T10:04:11-0500

Solution:

Ag+ + e- → Ag

Z = 1


Faraday's 1st law can be used:

n = (I × t) / (Z × F)

where:

n = number of moles of electrons (mol)

I = current (A)

t = total time the constant current was applied (s)

Z = number of electrons

F = Faraday constant, 96485 (A s mol-1)


Thus:

t = (n × Z × F) / (I)

t = (0.56 mol × 1 × 96485 A s mol-1) / (5.0 A) = 10806.32 s

t = (10806.32 s) × (1 min / 60 s) = 180 min

t = (10806.32 s) × (1 h / 3600 s) = 3 h


Answer: 3 hours or 180 min or 10806.32 s